OpenStudy (anonymous):

ann has the 5 newest music CD's which is 3 less than twice the amount that bob has.

OpenStudy (anonymous):

Let x = number of CDs Bob has. Ann is on the right side of the equation, Bob on the left. 2x - 3 = 5 Says that 2 times the number of CDs Bob has, less 3, is 5, the number that Ann has. Add 3 to both sides. 2x - 3 + 3 = 5 + 3 2x = 8 Divide both sides by 2. x = 4. Bob has 4 CDs.